# Anyon condensation in mixed-state topological order

### v3

@@ 1 Introduction and summary
In pure-state TOs, certain relations among them are provided by the well-known procedure, anyon condensation [13, 14, 15]. The procedure achieves the following. Pick a pure-state TO described by an MFC $\mathcal{B}_{1}$. A ‘nice’ anyon $A\in\mathcal{B}_{1}$ can be condensed. (Here, ‘nice’ anyons are given by connected étale algebras in $\mathcal{B}_{1}$ [16].) Condensing $A$, we obtain a new pure-state TO described by another MFC $\mathcal{B}_{2}$. (The two phases are separated by a wall described by $\mathcal{C}$ [^2] Mathematically, the new MFC $\mathcal{B}_{2}$ is given by the category $(\mathcal{B}_{1})_{A}^{0}$ of dyslectic $A$-modules. The theory $\mathcal{C}$ on the wall is given by the category $(\mathcal{B}_{1})_{A}$ of $A$-modules.) In this way, one can obtain a new pure-state TO from an old one. Following these developments, in this paper, we study anyon condensation in mixed-state TOs. [^3] As far as we know, anyon condensation in pre-MFC was first mentioned in the footnote 23 of [17] and studied more systematically in [18]. Namely, we pick a mixed-state TO described by a pre-MFC $\mathcal{B}_{1}$ and a condensable anyon $A\in\mathcal{B}_{1}$. Condensing $A$, we obtain a new phase described by $\mathcal{B}_{2}$. We identify $\mathcal{B}_{2}$ and the theory $\mathcal{C}$ on the wall separating them. (In the usual pure-state TOs, particles in $\mathcal{B}_{2}$ and $\mathcal{C}$ are called deconfined and confined, respectively.) As we comment below, anyon condensation in mixed-state TOs allows us to compute topological invariant of such phases suggested in [10].
@@ 1.1 Summary of main results
[…]
@@
Remark. We added the superscript “ph” to emphasize this is not a mathematical theorem, but a physical result. More precisely, it is mathematically known [19, 20, 21] that connected étale algebras in pre-MFCs can be condensed. However, to the best of our knowledge, it is not mathematically known whether all condensable anyons are exhausted by connected étale algebras or not. Our physical consideration claims this is the case. Unlike this result, the other three are mathematical results.
Interestingly, some condensations give MFCs $\mathcal{B}_{2}$ describing pure-state TOs. We clarified when this happens:
@@
Theorem. [26, 27] All mixed-state topological orders described by positive symmetric pre-modular fusion categories belong to the same equivalence class with $\mathcal{A}^{\text{min}}\simeq$$\text{Vect}_{\mathbb{C}}$.
@@ 1.2 Procedure and examples of anyon condensation
Having explained the procedure abstractly, let us see these in concrete examples. We study two simple examples: one from the familiar pure-state TO (1) Toric Code, and another from mixed-state TO (2) Rep$(S_{3})$. For the latter, mathematical anyon condensation was explained in [18].
@@
[…]
@@
Connected étale algebra $A$ $\mathcal{B}_{A}\simeq\mathcal{B}^{0}_{A}$ rank($\mathcal{B}_{A}$)
@@ 2 Condensable anyons in mixed-state topological order
$A$$A$$A$
Figure 4: The particle $A$ is trivial.
@@
$0\to X\to A\otimes_{\mathcal{B}_{1}}A\cong A\oplus X\stackrel{{\scriptstyle\mu}}{{\to}}A\to 0$
@@
$1\otimes_{\mathcal{B}_{1}}A$$A\otimes_{\mathcal{B}_{1}}A$$A$$A$$u\otimes_{\mathcal{B}_{1}}id_{A}$$l_{A}$$\mu$
@@
To see the unit axiom of an algebra $(A,\mu,u)$, pick a point inside a bubble. We view the point supports the trivial particle $A\in\mathcal{B}_{2}$. Thanks to (2.2), we can also view the particle as $A\otimes_{\mathcal{B}_{2}}A\in\mathcal{B}_{2}$. Now, as usual, rewind the tape until the point is uncovered by the bubble. It should be invariant whether we view the point now supports $A\in\mathcal{B}_{1}$ or $A\otimes_{\mathcal{B}_{1}}A\in\mathcal{B}_{1}$. Since we can further view $A\in\mathcal{B}_{1}$ as fusions $11\otimes_{\mathcal{B}_{1}}A\cong A\cong A\otimes_{\mathcal{B}_{1}}$ 1, we should get various equivalent viewpoints: 1) view $1\otimes_{\mathcal{B}_{1}}A$ as $A$, or 2) turn $1\otimes_{\mathcal{B}_{1}}A$ to $A\otimes_{\mathcal{B}_{1}}A$ via (2.2), and to $A$ via (2.3). The first operation is given by $id_{A}\cdot l_{A}$, and the second by $\mu\cdot(u\otimes_{\mathcal{B}_{1}}id_{A})$. Our inability to distinguish the two requires
@@
$A\otimes_{\mathcal{B}_{1}}1$$A\otimes_{\mathcal{B}_{1}}A$$A$$A$$id_{A}\otimes_{\mathcal{B}_{1}}u$$r_{A}$$\mu$
@@
$\quad\quad\quad m\otimes_{\mathcal{B}_{1}}A$$p$$m$
@@
$m$$A$$A$$m$$A$$m$$A$$m$$(id_{m}\otimes_{\mathcal{B}_{1}}\mu)\cdot\alpha^{\mathcal{B}_{1}}_{m,A,A}$$p\otimes_{\mathcal{B}_{1}}id_{A}$$p$$p$
@@
$m\otimes_{\mathcal{B}_{1}}1$$m\otimes_{\mathcal{B}_{1}}A$$m$$m$$id_{m}\otimes_{\mathcal{B}_{1}}u$$r_{m}$$p$
@@
$m\quad$$\quad A$$m\quad$$\quad A$
@@ 3.3 Example 3: $\mathcal{B}\simeq\text{Rep}(S_{4})$
[…]
@@
Since $Y$ is self-dual, consistency with the unique vacuum (1.3) demands $Y$ to split into two simple objects with quantum dimensions one
@@
$\mathcal{B}\stackrel{{\scriptstyle 1\oplus Y}}{{\longrightarrow}}\mathcal{B}^{\prime}_{A}\stackrel{{\scriptstyle 1\oplus X}}{{\longrightarrow}}\text{result}.$
@@ 3.4 Example 4: $\mathcal{B}\simeq\text{TY}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$
[…]

### v4

@@ 1 Introduction and summary
In pure-state TOs, certain relations among them are provided by the well-known procedure, anyon condensation [13, 14, 15]. The procedure achieves the following. Pick a pure-state TO described by an MFC $\mathcal{B}_{1}$. A ‘nice’ anyon $A\in\mathcal{B}_{1}$ can be condensed. (Here, ‘nice’ anyons are given by connected étale algebras in $\mathcal{B}_{1}$ [16].) Condensing $A$, we obtain a new pure-state TO described by another MFC $\mathcal{B}_{2}$. (The two phases are separated by a wall described by $\mathcal{C}$. [^2] Mathematically, the new MFC $\mathcal{B}_{2}$ is given by the category $(\mathcal{B}_{1})_{A}^{0}$ of dyslectic $A$-modules. The theory $\mathcal{C}$ on the wall is given by the category $(\mathcal{B}_{1})_{A}$ of $A$-modules.) In this way, one can obtain a new pure-state TO from an old one. Following these developments, in this paper, we study anyon condensation in mixed-state TOs. [^3] As far as we know, anyon condensation in pre-MFC was first mentioned in the footnote 23 of [17] and studied more systematically in [18]. Namely, we pick a mixed-state TO described by a pre-MFC $\mathcal{B}_{1}$ and a condensable anyon $A\in\mathcal{B}_{1}$. Condensing $A$, we obtain a new phase described by $\mathcal{B}_{2}$. We identify $\mathcal{B}_{2}$ and the theory $\mathcal{C}$ on the wall separating them. (In the usual pure-state TOs, particles in $\mathcal{B}_{2}$ and $\mathcal{C}$ are called deconfined and confined, respectively.) As we comment below, anyon condensation in mixed-state TOs allows us to compute topological invariant of such phases suggested in [10].
@@ 1.1 Summary of main results
Conjecture. Condensable anyons in a mixed-state topological order $\mathcal{B}_{1}$ are given by connected étale algebras $A\in\mathcal{B}_{1}$. They give the vacuum $A\cong 1_{\mathcal{B}_{2}}\in\mathcal{B}_{2}$ (denoted by 0 in later sections) in a new phase. The new phase is given by the category $(\mathcal{B}_{1})_{A}^{0}$ of dyslectic (right) $A$-modules. The two phases are separated by a wall given by the category $(\mathcal{B}_{1})_{A}$ of (right) $A$-modules.
@@
Remark. It is mathematically known [19, 20, 21] that connected étale algebras in pre-MFCs can be condensed. However, to the best of our knowledge, it is not mathematically known whether all condensable anyons are exhausted by connected étale algebras or not. Our physical consideration claims this is the case.
Interestingly, some condensations give MFCs $\mathcal{B}_{2}$ describing pure-state TOs. By finding mathematical results, we found when exactly this happens:
@@
[…]
@@ 1.2 Procedure and examples of anyon condensation
[…]
@@
Connected étale algebra AA $\mathcal{B}_{A}$ rank($\mathcal{B}_{A}$) $\mathcal{B}^{0}_{A}$ rank($\mathcal{B}^{0}_{A}$)
@@
Connected étale algebra AA $\mathcal{B}_{A}\simeq\mathcal{B}^{0}_{A}$ rank($\mathcal{B}_{A}$)
@@ 2 Condensable anyons in mixed-state topological order
AAAAAA
Figure 4: The particle AA is trivial.
@@
$0\to X\to A\otimes_{\mathcal{B}_{1}}A\cong A\oplus X\stackrel{{\scriptstyle\mu}}{{\to}}A\to 0.$
@@
$1\otimes_{\mathcal{B}_{1}}A$$\otimes_{\mathcal{B}_{1}}AAA$$Au\otimes_{\mathcal{B}_{1}}id_{A}$$l_{A}$$\mu$
@@
To see the unit axiom of an algebra $(A,\mu,u)$, pick a point inside a bubble. We view the point supports the trivial particle $A\in\mathcal{B}_{2}$. Thanks to (2.2), we can also view the particle as $A\otimes_{\mathcal{B}_{2}}A\in\mathcal{B}_{2}$. Now, as usual, rewind the tape until the point is uncovered by the bubble. It should be invariant whether we view the point now supports $A\in\mathcal{B}_{1}$ or $A\otimes_{\mathcal{B}_{1}}A\in\mathcal{B}_{1}$. Since we can further view $A\in\mathcal{B}_{1}$ as fusions $1\otimes_{\mathcal{B}_{1}}A\cong A\cong A\otimes_{\mathcal{B}_{1}}1$, we should get various equivalent viewpoints: 1) view $1\otimes_{\mathcal{B}_{1}}A$ as $A$, or 2) turn $1\otimes_{\mathcal{B}_{1}}A$ to $A\otimes_{\mathcal{B}_{1}}A$ via (2.2), and to $A$ via (2.3). The first operation is given by $id_{A}\cdot l_{A}$, and the second by $\mu\cdot(u\otimes_{\mathcal{B}_{1}}id_{A})$. Our inability to distinguish the two requires
@@
[…]
@@
$\quad\quad\quad m\otimes_{\mathcal{B}_{1}}A$ppmm
@@
[…]
@@
[…]
@@
$m\quadA\quad$ $m\quadA\quad A$
@@ 3.3 Example 3: $\mathcal{B}\simeq\text{Rep}(S_{4})$
$A\cong 1\oplus X,\ 1\oplus Y,\ 1\oplus X\oplus 2Y,\ 1\oplus X\oplus 2Y\oplus 3Z\oplus 3W.$
@@
Since the RHS contains the vacuum, we learn
$Y\cong Y^{*}.$
If $Y$ were simple, the fusion contradicts the uniqueness of the vacuum (1.3). Thus, $Y$ should split into two simple objects with quantum dimensions one
@@
$\mathcal{B}\stackrel{{\scriptstyle 1\oplus Y}}{{\longrightarrow}}\mathcal{B}_{A}\stackrel{{\scriptstyle 1\oplus X}}{{\longrightarrow}}\text{result}.$
@@ 3.4 Example 4: $\mathcal{B}\simeq\text{TY}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$
and conformal dimensions [^17] One can find these conformal dimensions noticing that the pre-MFC is a rank five braided fusion subcategory of $\text{Ising}\boxtimes\text{Ising}$ [30].